# A Survey of Safe Reinforcement Learning and Constrained MDPs: A Technical Survey on Single-Agent and Multi-Agent Safety

### v1

@@ 3.3 Constraint Types and Safety Specifications
Robotics — Speed vs. Safety Trade-off Achiam et al. (2017); Berkenkamp et al. (2017); Chow et al. (2018a): A mobile robot navigating in an environment may aim to maximize the reward associated with reaching the goal quickly. However, it also needs to minimize the probability of collisions and energy consumption. Here, speed contributes positively to the reward, while collisions and energy usage are treated as negative costs. The CMDP formulation could enforce a maximum acceptable collision rate and energy budget, leading to an explicit safety-performance trade-off.
Autonomous Driving — Travel Time vs. Accident Risk Dalal et al. (2018); Shalev-Shwartz et al. (2017); Zheng and Gu (2024): In autonomous driving, agents aim to minimize the expected travel time while simultaneously ensuring a low probability of accidents. The agent faces a trade-off between driving faster (leading to higher reward) and maintaining safe distances or reduced speeds to avoid collisions (cost). The Pareto frontier consists of policies ranging from conservative (low accident risk, long travel time) to aggressive (low travel time, high accident risk). SafeRL selects a policy on this frontier according to the safety constraint.
Energy Systems — Power Supply vs. Cost and Reliability Bitar and Low (2012); Dall’Anese et al. (2015): In power grid management, the agent may aim to optimize electricity production to meet demand (reward) while minimizing costs associated with fuel consumption and the risk of violating reliability standards (costs). This problem naturally involves multiple objectives: maximizing supply quality and minimizing operational risks and costs.
Healthcare — Treatment Success vs. Adverse Effects Gottesman et al. (2019); Raghu et al. (2017): In medical decision-making, an RL agent may need to maximize treatment efficacy while minimizing adverse effects or treatment toxicity. For example, maximizing patient recovery speed could conflict with the need to limit drug dosage to avoid harmful side effects. A CMDP constraint could limit the expected cumulative adverse effects to a tolerable threshold, enforcing safety.
Manufacturing — Production Efficiency vs. Maintenance Costs Chung et al. (2020); Khouja (2003): In automated manufacturing, increasing production speed or output (reward) may result in higher machine wear and maintenance costs (costs). A CMDP-based SafeRL framework may impose a constraint on expected maintenance cost or machine degradation, forcing the agent to balance throughput and longevity.
Drone Swarms — Task Completion vs. Communication Load Gu et al. (2023); Yang et al. (2020b): In multi-drone systems, agents may wish to maximize task completion rates (reward) while minimizing communication overhead (cost). Communication constraints can act as safety constraints in environments with bandwidth limitations or interference risks.
@@
Temporal logic specifications Alshiekh et al. (2018a); ElSayed-Aly et al. (2021); Turchetta et al. (2016); Wabersich and Zeilinger (2018): In some safety-critical settings, the safety requirement is given as a formal temporal logic formula (e.g., “always avoid region X unless Y happens”). Such logic specifications can be converted to automata and then to reward/cost functions or shields that enforce them Alshiekh et al. (2018a); ElSayed-Aly et al. (2021). While not a traditional CMDP constraint, they can often be incorporated by extending the state space to include automaton states representing the satisfaction of the formula. Specifically, in Alshiekh et al. (2018a), Alshiekh et al. (2018) introduced safe reinforcement learning via shielding; they used LTL (Linear Temporal Logic) specifications to construct shields for RL agents. In Wabersich and Zeilinger (2018) Wabersich and Zeilinger (2018), linear model predictive safety certification for learning-based control was employed. Although it focused on model predictive safety, their framework is capable of incorporating logic-based safety constraints. The paper ElSayed-Aly et al. (2021) extends shield-based safe RL to the multi-agent setting using temporal logic specifications. The paper Turchetta et al. (2016) while focused on safe exploration, their work shows how formal safety specifications can be integrated into exploration guarantees. One of the older but influential paper Sadigh et al. (2016) shows how specifications from temporal logic can shape safe planning.
@@ 3.4 Theoretical Results
Theorem 3.1 (Optimal Policy for CMDP Altman (1999)).
For a finite CMDP with bounded rewards and costs, there exists an optimal policy that attains the maximum in (1) (and corresponding optimal dual variables). Moreover, there exists an optimal policy that is stationary (time-independent) and can be chosen to be deterministic with respect to actions at all but possibly a measure-zero set of states. In practice, optimal policies may randomize between a small number of deterministic policies if needed to exactly satisfy constraints.
@@
Theorem 3.2 (Lagrange Duality Altman (1999)).
Under mild regularity conditions (e.g., finite state/action or convexity in policy space), The strong duality holds for the CMDP problem. That is,
and solving the dual yields the primal optimum. The optimal dual variables provide valuable information: if , then the -th constraint is active (tight) at the optimum; if , the optimum policy naturally satisfies -th constraint with some slack.
@@
Proposition 3.3 (Policy Gradient for Constrained Objectives).
If the policy is parameterized by (e.g., a neural network), one can derive gradients for the constrained problem. For instance, using the Lagrangian, the gradient of with respect to is
This leads to constrained policy gradient algorithms, where is updated in the direction of and is updated in the direction of . Many actor-critic style SafeRL methods employ this simultaneous gradient update (a form of primal-dual gradient descent) Chow et al. (2018b).
@@ 4.1 Lagrangian-based Policy Optimization
Lagrangian Actor-Critic: In this approach, one augments the standard RL loss with penalty terms for constraint costs. For example, one can define a penalized reward (eqn (11) in Tessler et al. (2019)) for a single-constraint problem, where is treated as a learnable parameter. An actor-critic algorithm (Algorithm 1 in Tessler et al. (2019)) can then be used: The actor (policy ) is updated with respect to the penalized objective , using policy gradient or other optimization. The critic(s) estimate both the value of the reward and the cost (often one critic for and one for ). The Lagrange multiplier is updated by gradient ascent on the constraint satisfaction term, e.g. . This simple scheme is often called the Lagrange method or reward shaping method in safe RL. It was used in early safe deep RL implementations (e.g., Tessler et al. (2019) for safe DQN with constraints, and policy-gradient variants, i.e., Trust Region Policy Optimization and Proximal Policy Optimization in Ray et al. (2019)). While straightforward, a drawback is that the penalty coefficient can be hard to tune (“Our baseline results for constrained RL indicate a need for stronger and/or better-tuned algorithms to succeed on Safety Gym environments” as quoted in Ray et al. (2019)) and the method does not guarantee strict constraint satisfaction until convergence. The agent might violate constraints during learning if is not large enough, or conversely, learn too slowly if is too large initially.
@@
Projected Lagrangian (Constrained Policy Optimization): Achiam et al. Achiam et al. (2017) introduced Constrained Policy Optimization (CPO), a landmark algorithm that improves upon the basic Lagrangian method by ensuring each policy update is safe. CPO is built on trust-region policy optimization: At each iteration, it solves a local constrained optimization: maximize policy improvement subject to a constraint that the cost does not increase beyond a small tolerance. This is done by a quadratic approximation of the objective and a linear approximation of the constraints (using policy gradient and cost gradient), then solving a convex subproblem. If the proposed update violates the constraint (predicted cost increase too high), CPO backtracks or projects the policy update to the nearest feasible update. CPO provides theoretical guarantees of near-constraint satisfaction at each iteration: essentially, it never overshoots the constraint by more than a certain second-order error term, keeping training safe. CPO demonstrated that one can train neural network policies for control tasks while maintaining safety throughout training Achiam et al. (2017). It was the first general-purpose safe RL algorithm with such guarantees. However, CPO is more complex and computationally heavier than standard policy gradient (due to solving the constrained optimization subproblem each step). It also requires a reliable estimation of the cost value and cost advantage, which can be challenging. Many subsequent works have built on or modified CPO: PCPO (Projection-based CPO): an algorithm that explicitly projects the policy gradient to the feasible set defined by constraint gradients Yang et al. (2020a). It is a simplification that avoids solving a quadratic program but still aims to keep updates safe by geometric projection.
@@
TRPO-Lagrangian: A simpler baseline where one applies a trust-region update on the penalized objective instead of solving a constrained QP. This does not guarantee strict feasibility but often empirically manages constraint violations by proper adaptation. OpenAI’s Safety Gym benchmark release OpenAI (2019); Ray et al. (2019) used such baselines111https://github.com/openai/safety-starter-agents.
@@
Actor-Critic with Lyapunov: Chow et al. Chow et al. (2018a) proposed using a Lyapunov function (a monotonic function of the cost-to-go) to derive a safe update rule. They ensure the new policy does not increase a Lyapunov function, which in turn guarantees the constraint remains satisfied. This can be seen as another form of trust-region or projection method specialized using Lyapunov theory.
@@
Off-policy and Model-based extensions: While most policy optimization methods are on-policy, there have been adaptations to off-policy learning: Safe DDPG or TD3: by incorporating a cost critic and Lagrange multiplier, one can train deterministic policies (as in DDPG) with a constraint. For example, a constrained variant of TD3 (Twin Delayed Deep Deterministic Policy GradientFujimoto et al. (2018)222https://spinningup.openai.com/en/latest/algorithms/td3.html) was proposed by Kostrikov et al. (2021); Lyu and Liu (2021).
@@ 4.2 Safety Shields and Action Correction
Another category of SafeRL methods focuses on safe exploration: how to prevent an agent from ever taking an action that could lead to catastrophe. These methods act as a layer on top of any standard RL algorithm, modifying or filtering its actions:
@@
Shielding via formal methods: Alshiekh et al. Alshiekh et al. (2018b) and later ElSayed-Aly et al. ElSayed-Aly et al. (2021) (extended to multiagents) use formal verification and temporal logic to build shields. The idea is to pre-compute a set of forbidden state-action pairs using model checking of an abstract model, or to synthesize a runtime observer from a formal specification. The shield then blocks any action that would lead into a bad state (violating the LTL safety specification) in finite steps. In multi-agent settings, as ElSayed-Aly et al. (2021) shows, one can have a centralized shield watching over joint actions or distributed shields for each agent.
@@ 4.3 Risk-Sensitive and Distributional Methods
Although our focus is on constraint-based SafeRL, a brief mention of risk-sensitive RL is warranted as an alternative approach: In risk-sensitive RL, instead of constraints, the optimization criterion itself is altered to account for risk. For example, one might maximize where is a concave utility (exponential utility gives risk-aversion) or maximize of return at some confidence level . Tamar et al. Tamar et al. (2015) and others have developed policy gradient methods for CVaR. These effectively try to ensure with high probability the return is above some level, which is conceptually similar to constraints on probabilities of bad events. Distributional RL (as popularized by Bellemare et al. (2017)) learns the full distribution of returns. One can combine distributional RL with safety by focusing on the lower tail of the return distribution to ensure it is above some threshold. This is another way to encode safety without explicit constraints. Risk-sensitive methods can sometimes be converted into CMDP style constraints. For instance, requiring CVaR is a constraint on a specific risk measure of cost. Solving such constraints often introduces auxiliary variables or uses sample-based approximations. While we do not detail these methods here, they are part of the broader SafeRL toolbox.
@@ 4.4 Safe Multi-Agent Reinforcement Learning (SafeMARL)
SafeMARL extends the ideas above to multi-agent systems Albrecht et al. (2024); Weiss (1999); Wooldridge (2009); Shoham and Leyton-Brown (2008); Weiss (1996). We consider environments with agents, indexed by . A convenient formal model is a constrained Markov game, defined by . Here each agent chooses an action , forming a joint action that causes state transitions via . Each agent can receive an individual reward and has possibly its own set of cost functions for .
@@
In competitive or general-sum SafeMARL, each agent has its own reward to maximize, and constraints might be individual (each agent has its own safety requirement) or shared (environment-level safety that everyone needs to uphold, like traffic rules). The solution concept might be a safe equilibrium Ganzfried (2022) (e.g., a Nash equilibrium that respects constraints, a related work Everitt et al. (2019) considers safe strategies for players) rather than a single policy optimization. Most existing work in SafeMARL addresses cooperative settings, since even standard MARL is most tractable in either fully cooperative (centralized training for a team) or fully competitive (two-player zero-sum) cases.
We highlight a few key approaches:
@@
Multi-agent Credit Assignment for Safety: In multi-agent RL, credit assignment (determining each agent’s contribution to global reward) is crucial. Similarly, for safety, one might need to assign “blame” or responsibility to individual agents for a safety violation. Approaches like difference rewards or shaped team rewards can be used to ensure each agent gets feedback about how its actions affected the global outcome. For SafeRL, one could design each agent’s cost signal such that it corresponds to the marginal increase in global risk due to that agent. Some initial works have considered approaches where each agent considering the safety constraints of others Gu et al. (2023), though a general solution is open research (we outline this as a problem later).
@@
Benchmarking SafeMARL: The progress in SafeMARL has been accelerated by the introduction of benchmarks. Gu et al. Gu et al. (2023) provided Safe Multi-Agent MuJoCo444https://github.com/chauncygu/Safe-Multi-Agent-Mujoco, Safe MARobosuite, and Safe MA-IsaacGym, which are multi-robot simulation tasks with safety constraints (like torque limits or collision constraints). These environments allow systematic evaluation of SafeMARL algorithms in settings requiring coordination. Similarly, for single-agent SafeRL, OpenAI’s Safety Gym555https://github.com/openai/safety-gym Ray et al. (2019) introduced a suite of continuous control tasks with hazards and constraints, which has become a standard testbed.
@@
| Lyapunov-based Policy Optimization Chow et al. (2018a) | Uses a Lyapunov function (cost critic) to constrain updates. Guarantees decrease in an upper bound of cost. |

### v2

@@ 3.3 Constraint Types and Safety Specifications
Robotics –Speed vs. Safety Trade-off Achiam et al. (2017); Berkenkamp et al. (2017); Chow et al. (2018): A mobile robot navigating in an environment may aim to maximize the reward associated with reaching the goal quickly. However, it also needs to minimize the probability of collisions and energy consumption. Here, speed contributes positively to the reward, while collisions and energy usage are treated as negative costs. The CMDP formulation could enforce a maximum acceptable collision rate and energy budget, leading to an explicit safety-performance trade-off.
Autonomous Driving –Travel Time vs. Accident Risk Dalal et al. (2018); Shalev-Shwartz et al. (2018); Zheng and Gu (2025); Nguyen and Han (2023): In autonomous driving, agents aim to minimize the expected travel time while simultaneously ensuring a low probability of accidents. The agent faces a trade-off between driving faster (leading to higher reward) and maintaining safe distances or reduced speeds to avoid collisions (cost). The Pareto frontier consists of policies ranging from conservative (low accident risk, long travel time) to aggressive (low travel time, high accident risk). SafeRL selects a policy on this frontier according to the safety constraint.
Energy Systems –Power Supply vs. Cost and Reliability Bitar and Xu (2017); Dall’Anese et al. (2017): In power grid management, the agent may aim to optimize electricity production to meet demand (reward) while minimizing costs associated with fuel consumption and the risk of violating reliability standards (costs). This problem naturally involves multiple objectives: maximizing supply quality and minimizing operational risks and costs.
Healthcare: Treatment Success vs. Adverse Effects Gottesman et al. (2019); Raghu et al. (2018): In medical decision-making, an RL agent may need to maximize treatment efficacy while minimizing adverse effects or treatment toxicity. For example, maximizing patient recovery speed could conflict with the need to limit drug dosage to avoid harmful side effects. A CMDP constraint could limit the expected cumulative adverse effects to a tolerable threshold, enforcing safety.
Manufacturing: Production Efficiency vs. Maintenance Costs Li et al. (2023); Siraskar et al. (2023); Chen and Zhou (2025): In automated manufacturing, increasing production speed or output (reward) may result in higher machine wear and maintenance costs (costs). A CMDP-based SafeRL framework may impose a constraint on expected maintenance cost or machine degradation, forcing the agent to balance throughput and longevity.
Drone Swarms: Task Completion vs. Communication Load Gu et al. (2023): In multi-drone systems, agents may wish to maximize task completion rates (reward) while minimizing communication overhead (cost). Communication constraints can act as safety constraints in environments with bandwidth limitations or interference risks.
@@
Temporal logic specifications Alshiekh et al. (2018); ElSayed-Aly et al. (2021); Turchetta et al. (2016); Wabersich and Zeilinger (2018): In some safety-critical settings, the safety requirement is given as a formal temporal logic formula (e.g., “always avoid region X unless Y happens”). Such logic specifications can be converted to automata and then to reward/cost functions or shields that enforce them Alshiekh et al. (2018); ElSayed-Aly et al. (2021). While not a traditional CMDP constraint, they can often be incorporated by extending the state space to include automaton states representing the satisfaction of the formula. Specifically, in Alshiekh et al. (2018), Alshiekh et al. (2018) introduced safe reinforcement learning via shielding; they used LTL (Linear Temporal Logic) specifications to construct shields for RL agents. In Wabersich and Zeilinger (2018) Wabersich and Zeilinger (2018), linear model predictive safety certification for learning-based control was employed. Although it focused on model predictive safety, their framework is capable of incorporating logic-based safety constraints. The paper ElSayed-Aly et al. (2021) extends shield-based safe RL to the multi-agent setting using temporal logic specifications. The paper Turchetta et al. (2016) while focused on safe exploration, their work shows how formal safety specifications can be integrated into exploration guarantees. One of the older but influential paper Sadigh et al. (2016) shows how specifications from temporal logic can shape safe planning.
@@ 3.4 Theoretical Results
Sketch of proof, detail in Achiam et al. (2017).
(1) Reward-shaping identity and surrogate. For any define and
Using the discounted visitation measures one shows
and by adding/subtracting and applying Hölder’s inequality, Achiam et al. (2017) obtain the two-sided bound
where .
(2) From to statewise TV between policies. Bound the shift in discounted visitation by the average per-state TV between action distributions, yielding
Substituting this into the previous display gives the Theorem 1 bounds in Achiam et al. (2017).
These bounds are significant as they justify using the expected advantage (the first term inside the expectation) as a surrogate objective for policy optimization. The bounds formally characterize the worst-case approximation error (the second term) that arises from using the state distribution of the old policy instead of the new policy’s distribution .
@@ 4.1 Lagrangian-based Policy Optimization
Lagrangian Actor-Critic: In this approach, one augments the standard RL loss with penalty terms for constraint costs. For example, one can define a penalized reward (eqn (11) in Tessler et al. (2019)) for a single-constraint problem, where is treated as a learnable parameter. An actor-critic algorithm (Algorithm 1 in Tessler et al. (2019)) can then be used: The actor (policy ) is updated with respect to the penalized objective , using policy gradient or other optimization. The critic(s) estimate both the value of the reward and the cost (often one critic for and one for ). The Lagrange multiplier is updated by gradient ascent on the constraint satisfaction term, e.g. . This simple scheme is often called the Lagrange method or reward shaping method in safe RL. It was used in early safe deep RL implementations (e.g., Tessler et al. (2019) for safe DQN with constraints, and policy-gradient variants, i.e., Trust Region Policy Optimization and Proximal Policy Optimization in Ray et al. (2019a)). While straightforward, a drawback is that the penalty coefficient can be hard to tune (“Our baseline results for constrained RL indicate a need for stronger and/or better-tuned algorithms to succeed on Safety Gym environments” as quoted in Ray et al. (2019a)) and the method does not guarantee strict constraint satisfaction until convergence. The agent might violate constraints during learning if is not large enough, or conversely, learn too slowly if is too large initially.
@@
Projected Lagrangian (Constrained Policy Optimization): Achiam et al. Achiam et al. (2017) introduced Constrained Policy Optimization (CPO), a landmark algorithm that improves upon the basic Lagrangian method by ensuring each policy update is safe. CPO is built on trust-region policy optimization: At each iteration, it solves a local constrained optimization: maximize policy improvement subject to a constraint that the cost does not increase beyond a small tolerance. This is done by a quadratic approximation of the objective and a linear approximation of the constraints (using policy gradient and cost gradient), then solving a convex subproblem. If the proposed update violates the constraint (predicted cost increase too high), CPO backtracks or projects the policy update to the nearest feasible update. CPO provides theoretical guarantees of near-constraint satisfaction at each iteration: essentially, it never overshoots the constraint by more than a certain second-order error term, keeping training safe. CPO demonstrated that one can train neural network policies for control tasks while maintaining safety throughout training Achiam et al. (2017). It was the first general-purpose safe RL algorithm with such guarantees. However, CPO is more complex and computationally heavier than standard policy gradient (due to solving the constrained optimization subproblem each step). It also requires a reliable estimation of the cost value and cost advantage, which can be challenging; for complex environments, cost estimates may be derived from a separate classifier trained to identify safe/unsafe actions Chirra et al. (2025).
Sketch of proof, detail in Achiam et al. (2017).
(1) Start from the cost performance bound. For any cost and policies , Achiam et al. (2017) give (Corollary 2):
where .
(2) Replace TV by average KL under a trust region. By Pinsker’s inequality and Jensen’s inequality,
If is produced by the CPO update with trust region , then
(3) Apply to the constrained subproblem and re-arrange. In the CPO subproblem, the surrogate constraint enforces Plugging and into the inequality of Step (1), and then substituting the TVKL bound from Step (2) yields
which is exactly the stated bound. ∎
Many subsequent works have built on or modified CPO: PCPO (Projection-based CPO): an algorithm that explicitly projects the policy gradient to the feasible set defined by constraint gradients Yang et al. (2020). It is a simplification that avoids solving a quadratic program but still aims to keep updates safe by geometric projection.
@@
TRPO-Lagrangian: A simpler baseline where one applies a trust-region update on the penalized objective instead of solving a constrained QP. This does not guarantee strict feasibility but often empirically manages constraint violations by proper adaptation. OpenAI’s Safety Gym benchmark release Ray et al. (2019b, a) used such baselines111https://github.com/openai/safety-starter-agents.
@@
Actor-Critic with Lyapunov: Chow et al. Chow et al. (2018) proposed using a Lyapunov function (a monotonic function of the cost-to-go) to derive a safe update rule. They ensure the new policy does not increase a Lyapunov function, which in turn guarantees the constraint remains satisfied. This can be seen as another form of trust-region or projection method specialized using Lyapunov theory.
@@
Off-policy and Model-based extensions: While most policy optimization methods are on-policy, there have been adaptations to off-policy learning: Safe DDPG or TD3: by incorporating a cost critic and Lagrange multiplier, one can train deterministic policies (as in DDPG) with a constraint. For example, a constrained variant of TD3 (Twin Delayed Deep Deterministic Policy GradientFujimoto et al. (2018)222https://spinningup.openai.com/en/latest/algorithms/td3.html) was proposed by Zhang et al. (2023); Yang et al. (2025).
@@ 4.2 Safety Shields and Action Correction
Another category of SafeRL methods focuses on safe exploration: how to prevent an agent from ever taking an action that could lead to catastrophe. These methods act as a layer on top of any standard RL algorithm, modifying or filtering its actions (see Figure 12):
@@
Shielding via formal methods: Alshiekh et al. Alshiekh et al. (2018) and later ElSayed-Aly et al. ElSayed-Aly et al. (2021) (extended to multiagents) use formal verification and temporal logic to build shields. The idea is to pre-compute a set of forbidden state-action pairs using model checking of an abstract model, or to synthesize a runtime observer from a formal specification. The shield then blocks any action that would lead into a bad state (violating the LTL safety specification) in finite steps. In multi-agent settings, as ElSayed-Aly et al. (2021) shows, one can have a centralized shield watching over joint actions or distributed shields for each agent.
@@ 4.3 Risk-Sensitive and Distributional Methods
Although our focus is on constraint-based SafeRL, a brief mention of risk-sensitive RL is warranted as an alternative approach: In risk-sensitive RL, instead of constraints, the optimization criterion itself is altered to account for risk. For example, one might maximize where is a concave utility (exponential utility gives risk-aversion) or maximize of return at some confidence level . Tamar et al. Tamar et al. (2015) and others have developed policy gradient methods for CVaR (see Bäuerle and Jaśkiewicz (2024) for an overview of risk-sensitive criteria in MDPs). These effectively try to ensure with high probability the return is above some level, which is conceptually similar to constraints on probabilities of bad events. Distributional RL (as popularized by Bellemare et al. (2017)) learns the full distribution of returns. One can combine distributional RL with safety by focusing on the lower tail of the return distribution to ensure it is above some threshold. This is another way to encode safety without explicit constraints. Risk-sensitive methods can sometimes be converted into CMDP style constraints. For instance, requiring CVaR is a constraint on a specific risk measure of cost. Solving such constraints often introduces auxiliary variables or uses sample-based approximations. While we do not detail these methods here, they are part of the broader SafeRL toolbox.
@@ 4.4 Safe Multi-Agent Reinforcement Learning (SafeMARL)
SafeMARL extends the ideas above to multi-agent systems Albrecht et al. (2024); Weiss (2000); Wooldridge (2009); Shoham and Leyton-Brown (2008). We consider environments with agents, indexed by . A convenient formal model is a constrained Markov game, defined by . Here each agent chooses an action , forming a joint action that causes state transitions via . Each agent can receive an individual reward and has possibly its own set of cost functions for .
@@
In competitive or general-sum SafeMARL, each agent has its own reward to maximize, and constraints might be individual (each agent has its own safety requirement) or shared (environment-level safety that everyone needs to uphold, like traffic rules). The solution concept might be a safe equilibrium Ganzfried (2023) (e.g., a Nash equilibrium that respects constraints, related works Altman and Shwartz (2000); Mccracken and Bowling (2004) consider constrained Markov games and safe strategies for players) rather than a single policy optimization. Most existing work in SafeMARL addresses cooperative settings, since even standard MARL is most tractable in either fully cooperative (centralized training for a team) or fully competitive (two-player zero-sum) cases.
We highlight a few key approaches (see Figure 13 for an architectural comparison):
Sketch of proof.
Step 1: Multi-agent surrogate decomposition. Let be the joint advantage under the current joint policy . By the multi-agent advantage decomposition (Lemma 1 in Gu et al. (2023)), for any ordering ,
Define the one-agent surrogate
A standard TRPO bound (applied in the multi-agent setting) gives
for .
Step 2: Sequential argmax and the identity at the old policy. By construction, . Each agent update is chosen to maximize the penalized surrogate , so replacing by yields a lower value. Summing over gives , proving monotonic improvement.
Step 3: Per-iteration feasibility via cost surrogates + KL budgets. For each agent and cost index , Lemma 2 in Gu et al. (2023) bounds the change of cost under a joint update:
with . Choosing agent-wise KL radii ensures that, starting from a feasible , the sequential update of keeps every other agent’s constraint bounded by its threshold; hence for all . This proves per-iteration constraint satisfaction.
Combining Steps 1–3 yields the proposition. ∎
@@
Multi-agent Credit Assignment for Safety: In multi-agent RL, credit assignment (determining each agent’s contribution to global reward) is crucial. Similarly, for safety, one might need to assign “blame” or responsibility to individual agents for a safety violation. Approaches like difference rewards Wolpert and Tumer (2001); Tumer and Wolpert (2004) or shaped team rewards Devlin and Kudenko (2011) can be used to ensure each agent gets feedback about how its actions affected the global outcome. For SafeRL, one could design each agent’s cost signal such that it corresponds to the marginal increase in global risk due to that agent. Some initial works have considered approaches where each agent considering the safety constraints of others Gu et al. (2023), though a general solution is open research (we outline this as a problem later).
@@
Benchmarking SafeMARL: The progress in SafeMARL has been accelerated by the introduction of benchmarks. Gu et al. Gu et al. (2023) provided Safe Multi-Agent MuJoCo444https://github.com/chauncygu/Safe-Multi-Agent-Mujoco, Safe MARobosuite, and Safe MA-IsaacGym, which are multi-robot simulation tasks with safety constraints (like torque limits or collision constraints). These environments allow systematic evaluation of SafeMARL algorithms in settings requiring coordination. Similarly, for single-agent SafeRL, OpenAI’s Safety Gym555https://github.com/openai/safety-gym Ray et al. (2019a) introduced a suite of continuous control tasks with hazards and constraints, which has become a standard testbed. More recently, Safety Gymnasium Ji et al. (2023a) provides a modernized successor with Gymnasium API support, and algorithm libraries such as OmniSafe Ji et al. (2023b) and Safe Policy Optimization (SafePO) Ji et al. (2023a) offer unified implementations of constrained RL baselines. For safe control benchmarking, Safe-Control-Gym Yuan et al. (2022) integrates constraint-aware control tasks for robotics. Multi-agent driving environments such as SMARTS Zhou et al. (2020) and single-agent driving suites like Highway-env Leurent (2018) further expand the available testbeds for safe RL research.
@@
| Lyapunov-based Policy Optimization Chow et al. (2018) | Uses a Lyapunov function (cost critic) to constrain updates. Guarantees decrease in an upper bound of cost. |
@@
5 Safe RL and Safe MARL Libraries
In Tables 2 and 3, we show various available popular libraries and environments for further research in safe RL and safe MARL.
Taxonomy of Safe Reinforcement Learning Libraries and Environments.
To consolidate the growing ecosystem of reproducible and standardized implementations in Safe Reinforcement Learning (SafeRL) and Safe Multi-Agent Reinforcement Learning (SafeMARL), Tables 2 and 3 summarize the major algorithmic libraries and environment suites currently used by the research community. Table 2 lists the most widely adopted open-source frameworks that implement constrained or risk-aware reinforcement learning algorithms under the Constrained Markov Decision Process (CMDP) formulation. These include high-quality benchmark suites such as OmniSafe, Safe Policy Optimization (SafePO), Safety Starter Agents, and Safe-Control-Gym. Each library is characterized by its category, algorithmic focus, GitHub repository, and the environments it supports. The table also provides typical application domains ranging from robotics and navigation to multi-agent coordination and associated references, thereby serving as a guide to reproducible SafeRL experimentation and comparison across methods.
Benchmarking Environments and Datasets.
Complementing the algorithmic libraries, Table 3 categorizes the principal benchmark environments and datasets that provide structured safety signals for policy learning. These include the Safety-Gymnasium and the legacy OpenAI Safety Gym for constrained navigation and manipulation, as well as the AI Safety Gridworlds and SafeLife platforms that test specification robustness and side-effect avoidance in discrete domains. More complex continuous and multi-agent settings are covered by simulation frameworks such as SMARTS for autonomous driving and Highway-env for risk-sensitive control. Each environment entry specifies whether it supports single-agent or multi-agent training, the structure of safety constraints or cost signals, representative tasks, and intended application areas. Together, these two tables form a comprehensive taxonomy of tools that underpin experimental research in safe and trustworthy reinforcement learning.
| Name | Category | Algorithms / Capabilities | Application Areas |
| OmniSafe Ji et al. (2023b) | Algorithm Library (Safe RL) | PPO/TRPO-Lagrangian, CPO, NPG-Lag, SAC-Lag; experiment management | Robotics, safe navigation |
| SafePO Ji et al. (2023a) | Safe RL & Safe MARL | Lagrangian & CPO-style variants; MARL support; training pipelines | Multi-agent coordination, safe control |
| Safety Starter Agents Ray et al. (2019a) | Baselines / Reference | CPO, PPO-Lag, TRPO-Lag, DDPG-Lag baselines | Benchmarking baselines |
| Safe-Control-Gym Yuan et al. (2022) | Benchmarking Suite | Constraint-aware control tasks; integrates SB3 and RL agents | Safe robotics & continuous control |
| Name | Agent Type | Key Features / Signals | Example Tasks | Citations |
| Safety Gymnasium | Both | Cost signals & constraints (hazards); Gymnasium API | Goal/Button/Push with hazards | Ji et al. (2023a) |
| OpenAI Safety Gym | Single | Original cost/constraint tasks (archived) | Point/Car/Doggo: Goal/Button/Push | Ray et al. (2019a) |
| AI Safety Gridworlds | Single | Side-effect & reward specification tests | Multiple gridworlds for specification failures | Leike et al. (2017) |
| SafeLife | Single | Procedurally generated levels; side-effect metrics | Life-like tasks minimizing side effects | Wainwright and Eckersley (2021) |
| SMARTS | Multi | Traffic simulation, risk modeling for multi-agent driving | Merging, intersection, adversarial driving | Zhou et al. (2020) |
| Highway-env | Single | Collision-avoidance objectives; Gym-compatible | Highway, merge, roundabout, parking | Leurent (2018) |